OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2). y = two thirdsx − 10 y = two thirdsx + 10 y = −three halvesx − 10 y = −three halvesx + 10

OpenStudy (anonymous):

first change 2x-3y = 3 into this form y = mx + b where m is the slope and b is the y-intercept perpendicular lines have a negative reciprocal slope of the other (-1/m) change the equation into the form y = mx + b to find m then find the equation for the perpendicular line

OpenStudy (anonymous):

i cant find it :/

OpenStudy (anonymous):

2x-3y=3 minus 2x from both sides -3y = 3 - 2x divide by -3 both sides y = (2/3)x - 1 so the slope of the line is 2/3 the slope of the perpendicular line will be -3/2 y = -(3/2)x + b substitute x = -8, y = 2 to find b

OpenStudy (anonymous):

-3y = -2x+3... divide -3 from y you get.. y=2/3x+1.. take your point.. and put it in point slope formula. y-y1 = 2/3(x+8).. so you get y-2 = 2/3(x +8) add 2 to both sides.. it is y=2/3x+10

OpenStudy (anonymous):

@cvickery perpendicular line has negative reciprocal (-1/m) slope, no? o,o

OpenStudy (anonymous):

Yeah you are correct on that @nphuongsun93 I derped.

OpenStudy (anonymous):

so0 hard :/

OpenStudy (anonymous):

lol @pleasehelpme54 do this anyway y = -(3/2)x + b substitute x = -8, y = 2 to find b

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

I believe you are asking someone to do all the work for you... Which I understand but he can't set it up for you any more clearer. 2 = -(3/2)(-8) + b.. 2 = 12 + b b = -10.. so the equation is.. y = (-3/2)x + - 10

OpenStudy (anonymous):

- 10* so its D on your thing.. I guess?

OpenStudy (anonymous):

y = -(3/2)x + b substitute x = -8, y = 2 to find b 2 = -(3/2)(-8) + b 2 = 12 + b b = -10 y = -(3/2)x - 10 it's C lol

OpenStudy (anonymous):

C yeah.. okay.. jeeze I am doing bad today with multiple choice.

OpenStudy (anonymous):

lol thnks